OpenStudy (anonymous):

Will medal for help :) Addition and subtraction equations <33 1/3x + 1/3(2x - 15) = 3 1/2

OpenStudy (anonymous):

go to mathway.com

OpenStudy (anonymous):

@dmex it gave me an incorrect answer :/

OpenStudy (anonymous):

did it give you \[\frac{ 17 }{ 2}\]

OpenStudy (anonymous):

@kewlkatcutie did it give you \[\frac{ 17 }{ 2 }\] or \[3\frac{ 1 }{ 2 }\]